OpenStudy (anonymous):

you are estimating the total distance of a road trip by measuring the distance on a map.on the map the distance between towns measure 3/12 inch,4/36 inch,and 8/18 inch,respectively.What is the total distance of the trip on the map.

OpenStudy (anonymous):

@mary.rojas help me

OpenStudy (anonymous):

@MathsStar help me

OpenStudy (anonymous):

@Euler271

OpenStudy (anonymous):

\[\frac{ 3 }{ 12 } +\frac{ 4 }{ 36 }+\frac{ 8 }{ 18 }\] to add fractions they must have the same denominator. we will force the same denominator by multiplying by the required factor. notice we are basically multiplying by one [x/x = 1] we will use the common multiple 36: \[\frac{ 3 }{ 12 }*\frac{ 3 }{ 3 } +\frac{ 4 }{ 36 }+\frac{ 8 }{ 18 }*\frac{ 2 }{ 2 }\]\[\frac{ 9 }{ 36 } + \frac{ 4 }{ 36 } + \frac{ 16 }{ 36 } = \frac{ 9+4+16 }{ 36 } = \frac{ 29 }{ 36 }\]

OpenStudy (anonymous):

thank you :)